OpenStudy (anonymous):

What are the solutions of the quadratic equation below? 2x2 - 2x - 9 = 0? it has to be a fraction please help.

OpenStudy (wolfboy):

I dont remember this right now but this may help. http://www.wolframalpha.com/input/?i=quadratic+answers+in+fractions+2x2++-++2x++-++9++%3D++0%3F+

OpenStudy (campbell_st):

well you need to general quadratic formula so the roots are real and irrational... after checking by using the discriminant

OpenStudy (unklerhaukus):

A Quadratic Equation\[ax^2+bx+c=0\] The Quadratic Formula\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Start by identifying the coefficients \(a\), \(b\), \(c\) Compare \[2x^2 - 2x - 9 = 0\]With \[ax^2+bx+c=0\] \[a=?\\ b=?\\ c=?\]

OpenStudy (unklerhaukus):

Can you identify the coefficients @Soccergirl52 ?

OpenStudy (anonymous):

i cannot :C its too hard >.<

OpenStudy (unklerhaukus):

Compare \[\color{orange}2x^2+\color{salmon}{−2}x+\color{cornflowerblue}{−9}=0\]With \[\color{orange}ax^2+\color{salmon}bx+\color{cornflowerblue}c=0\] \[\color{orange}a=?\\\color{salmon}b=?\\\color{cornflowerblue}c=?\]

OpenStudy (anonymous):

let me see..

OpenStudy (anonymous):

its hard >.< :*(

OpenStudy (anonymous):

x=(-b±√b^2-4ac)/2a x=(2±√4+72) /4 x=(1±√19)/2

OpenStudy (anonymous):

UnkleRhaukus has explained everything required...... First learn the basics like finding the coefficients..... Then it is very easy.... :)